Ann is now 25 years old and she is planning to start saving for retirement. She expects her income of $60,000 in the coming year to grow at the (nominal) rate of 5% a year until she retires at the age of 65. She wants to save a fixedpercentage of her income per year. She wants to save enough money to be able to consume per year 50% of her income (in real terms) just before retirement (at age 65) for 20 years. Assume the inflation rate of 3%, and a nominal rate of return on Ann’s savings of 6%. What fraction of income should Ann be saving?
Answer #1
Years to retirement 39 (65-26)
Income growth rate=5% 0.05
Current annual income $60,000
Annual Income just before retirement $402,285 (60000*(1.05^39)
Amount required per year during retirement $201,143 (402385*0.5)
Number of years during retirement 20
Nominal rate of return 6%
Inflation Rate 3%
Real Rate of return=R
(1+R)=1.06/1.03= 1.029126214
Real Rate of return=R= 0.029126214
Real Rate of return=R= 2.91%
Value of total retirement amount required at age 65 $3,016,812 (Using PV function of excel with Rate=2.91%, Nper=20,Pmt=-211200)
Total Savings Required in nominal dollar $3,016,812
Future value of Saving after 40 years=(Savings )*(1.06^(40-N))
N=Year of Saving
If 100% of Income are saved, the totalsavings:
Income in year 1(from today) $6,000
Income in year 2 (60000*1.05)
Income in year (N+1)=1.05*(Income in year (N))
N A B=A*(1.06^(40-N))
Year Saving (100%) Future Value of savings
1 $60,000 $582,210.45
2 $63,000 $576,717.90
3 $66,150 $571,277.16
4 $69,458 $565,887.76
5 $72,930 $560,549.19
6 $76,577 $555,260.99
7 $80,406 $550,022.68
8 $84,426 $544,833.79
9 $88,647 $539,693.85
10 $93,080 $534,602.39
11 $97,734 $529,558.98
12 $102,620 $524,563.14
13 $107,751 $519,614.43
14 $113,139 $514,712.40
15 $118,796 $509,856.63
16 $124,736 $505,046.66
17 $130,972 $500,282.07
18 $137,521 $495,562.43
19 $144,397 $490,887.31
20 $151,617 $486,256.30
21 $159,198 $481,668.97
22 $167,158 $477,124.93
23 $175,516 $472,623.75
24 $184,291 $468,165.03
25 $193,506 $463,748.38
26 $203,181 $459,373.40
27 $213,340 $455,039.69
28 $224,007 $450,746.86
29 $235,208 $446,494.53
30 $246,968 $442,282.32
31 $259,317 $438,109.84
32 $272,282 $433,976.73
33 $285,896 $429,882.61
34 $300,191 $425,827.11
35 $315,201 $421,809.88
36 $330,961 $417,830.54
37 $347,509 $413,888.74
38 $364,884 $409,984.13
39 $383,129 $406,116.36
40 $402,285 $402,285.07
SUM $19,474,375.35
Total Future value at 100% Savings $19,474,375
Future Value Required at age 65 $3,016,812
Savings Required 0.15491187 (3016812/19474375)
Percentage of Savings Required 15.49%